# MHE under parametric uncertainty - Robust state estimation without informative data

### v1

@@ 2.3 Standard MHE for State and Parameter Estimation
In the following, we introduce a standard MHE formulation to jointly estimate states and parameters, which is widely applied, see, e.g., [46, 5, 6, 7, 8, 9, 10, 12, 11, 13, 15, 14]. The MHE approach at time step $t$ considers past input data $\{u_{j}\}_{j=t-M_{t}}^{t-1}$ and estimates of past noises and outputs $\{\bar{w}_{j},\bar{y}_{j}\}_{j=t-M_{t}}^{t-1}$ within a window of length $M_{t}=\min\{t,M\}$, with horizon $M\in\mathbb{I}_{\geq 0}$, and the past state and parameter estimates, $\hat{x}_{t-M_{t}}$ and $\hat{\theta}_{t-M_{t}}$, respectively. Thereby, the MHE scheme optimizes over a sequence of $M_{t}$ disturbance estimates $\hat{w}_{\cdot|t}=\{\hat{w}_{j|t}\}_{j=t-M_{t}}^{t-1}$, and the initial state and the parameter estimates, $\hat{x}_{t-M_{t}|t}$ and $\hat{\theta}_{t}$, respectively. Using the system model (2.1), the disturbance sequence, initial state estimate, and parameter estimate define a sequence of state $\{\hat{x}_{j|t}\}_{j=t-M_{t}}^{t}$ and output $\{\hat{y}_{j|t}\}_{j=t-M_{t}}^{t-1}$ estimates. The cost function is chosen as
@@
We denote a (not necessarily unique) minimizer of (2.3) as $\hat{x}_{t-M_{t}|t}^{*}$, $*\hat{w}_{\cdot|t}^{*}$, and $\hat{\theta}_{t}^{*}$, with the corresponding state and output sequences denoted as $*\hat{x}_{\cdot|t}^{*}$ and $*\hat{y}_{\cdot|t}^{*}$, respectively. The resulting state and parameter estimates at time step $t$ are denoted as $\hat{x}_{t}=\hat{x}_{t|t}^{*}$ and $\hat{\theta}_{t}=\hat{\theta}_{t}^{*}$, respectively. The MHE scheme (2.3) is applied in receding horizon fashion. Thereby, the current state and parameter estimates are obtained at each time step $t$ by solving the (finite-horizon) MHE problem (2.3) based on the $M_{t}$ most recent inputs, noise and output estimates, and the prior weightings based on the state and parameter estimates at time step $t-M_{t}$.
Stability of such an MHE scheme can be established using tools from [40] if the augmented state $\begin{bmatrix}x_{t}^{\top}&\theta^{\top}\end{bmatrix}^{\top}$ is detectable, the horizon length $M$ is chosen sufficiently large, and if $\eta\in[0,1)$. In [47] it was shown that the underlying detectability conditions for the augmented state holds under a PE condition.
@@ 3 Motivating Example
Figure 1: Academic example (Section 3): Absolute state and parameter estimation errors resulting from a standard MHE approach for state and parameter estimation (2.3).
In this section, we provide an academic example to demonstrate that standard MHE formulations (Section 2.3) lack fundamental robust stability properties in the absence of PE. The presented example is inspired by [53, Sec. V], which demonstrates parameter drift in adaptive filtering. Consider the system
@@
$y_{t}=$ $\begin{bmatrix}x_{t}+[w_{t}]_{2}\\ \theta x_{t}+[w_{t}]_{3}\end{bmatrix},$ (16d)
with $a=0.99$, $[w_{t}]_{1}=0$, $[w_{t}]_{2}=-0.1$, and $[w_{t}]_{3}=0.1$ for all $t\in\mathbb{I}_{\geq 0}$. Thereby, the constant values of $[w_{t}]_{2}$ and $[w_{t}]_{3}$ can be interpreted as sensor biases, which often occur in practice. Further consider $x_{0}=1$, $\bar{x}_{0}=1$, $\theta=1$, and $\bar{\theta}_{0}=1.1$, and the MHE formulation (2.3) with $M=20$, $\eta=0.99$, $Q=I_{3}$, $R=I_{2}$, $P_{\mathrm{x}}=10$, $P_{\theta}=10$, and $\bar{w}_{t}=0$ and $\bar{y}_{t}=y_{t}$ for all $t\in\mathbb{I}_{\geq 0}$. Note that
@@
The input is zero over long periods of time and the state asymptotically converges to zero. Hence, the ratio of the excitation signal compared to the noise becomes very small. As shown in Figure 1, the parameter diverges for such a standard MHE formulation. Consequently, whenever the output contains informative data (when the input is equal to $1$), the drifted parameter estimate leads to an increasing state estimation error. This phenomena is known as bursting in classical adaptive control [54, 55] and is depicted in the zoom-in in Figure 1. Thus, robust convergence of the resulting state and parameter estimate is not guaranteed. Note that this problem equally occurs if an MHE scheme (2.3) without discounting ($\eta=1$) is applied. In the remainder of the paper, we propose a novel MHE formulation which overcomes this issue by using a different prior weighting, which allows us to establish stability in the state estimates even in the absence of PE.
@@ 4.1 Moving Horizon Estimation Formulation
Equivalently to the standard MHE formulation stated in Section 2.3, our proposed MHE at time step $t$ considers past input data $\{u_{j}\}_{j=t-M_{t}}^{t-1}$ and estimates of past disturbances and outputs $\{\bar{w}_{j},\bar{y}_{j}\}_{j=t-M_{t}}^{t-1}$ in a window of length $M_{t}=\min\{t,M\}$, with horizon $M\in\mathbb{I}_{\geq 0}$, and the past state estimate $\hat{x}_{t-M_{t}}$. However, instead of using the past parameter estimate $\hat{\theta}_{t-M_{t}}$, the proposed MHE formulation relies on the prior $\bar{\theta}_{0}$ of the unknown parameter. The objective of the optimization-based state estimation problem is chosen as
@@
$\min_{\hat{x}_{t-M_{t}|t},\hat{w}_{\cdot|t},\hat{\theta}_{t}}$ $V_{\mathrm{MHE}}^{x}(\hat{x}_{t-M_{t}|t},\hat{w}_{\cdot|t},\hat{y}_{\cdot|t},\hat{\theta}_{t},t)$ (18a)
@@
We denote a (non-unique) minimizer [^3] We assume that a minimizer to (2.3) always exists. Existence of such a minimizer can be ensured under mild assumptions, e.g., if $P_{2,\mathrm{x}},P_{2,\theta},Q,R\succ 0$, or the sets $\mathbb{W},\mathbb{Y},\mathbb{X},\Theta$ are compact, cf. [39, App. A.11]. of (4.1) as $\hat{x}_{t-M_{t}|t}^{*}$, $*\hat{w}_{\cdot|t}^{*}$, and $\hat{\theta}_{t}^{*}$, with the corresponding state and output sequences denoted as $*\hat{x}_{\cdot|t}^{*}$ and $*\hat{y}_{\cdot|t}^{*}$, respectively. The resulting state estimate at time step $t$ is denoted as
@@
The MHE estimator (4.1) is then applied in a receding horizon fashion. Thereby, the state estimate (19) at each time step $t$ is obtained by solving the optimization problem (4.1) based on the $M_{t}$ most recent inputs, noise and output estimates, and the prior weighting based on the state estimate at time step $t-M_{t}$.
@@ 4.2 Theoretical Analysis
In the following, we establish that the MHE (4.1) is an exponentially $\delta$-IOpS estimator according to Definition 2 provided the horizon $M$ is chosen larger than some lower bound.
@@
with $\rho\in[0,1)$. Then, the MHE estimator (4.1) is exponentially $\delta$-IOpS according to Definition 2, i.e, (9) holds for all $t\in\mathbb{I}_{\geq 0}$ with (uniform) constants $C_{1},\ C_{2},\ C_{3}>0$, and
ϵ=italic-$\epsilon=$ $\ \frac{4}{\sqrt{1-\rho}}\sqrt{\frac{\lambda_{\max}(P_{2,\theta})}{\lambda_{\min}(P_{1})}}\sqrt{\rho}^{M}\|\bar{\theta}_{0}-\theta\|,$ (22)
@@
The proof of Theorem 4.2 can be found in Appendix .1. Note that the bound (9) with (22) resulting from Theorem 4.2 implies that for $M$ going to infinity, the MHE (4.1) is an exponentially $\delta$-IOS estimator according to Definition 2. It follows that FIE, i.e., Problem (4.1) with $M_{t}=t$, is exponentially $\delta$-IOS.
@@
Let Assumption 1 hold. Then, the FIE estimator (4.1) with $M_{t}=t$ is exponentially $\delta$-IOS according to Definition 2.
Proof 4.4.
@@
The above corollary shows that the results from [50, Thm. 9] are recovered, compare also our discussion regarding functional estimation in Section 5.4 below.
@@
In case the true parameter is known, i.e., $\bar{\theta}_{0}=\theta$, it follows directly from Theorem 4.2, and in particular (22), that the estimation error (20) resulting from the proposed MHE (4.1) is robustly globally exponentially stable (with $\epsilon=0$) according to [56, Def. 1].
@@
Choosing $P_{2,\mathrm{x}}$, $P_{2,\theta}$, $R$, and $Q$ in (17) based on a $\delta$-IOOS Lyapunov function $W_{\delta}$ (Definition 3) might seem to limit the applicability of the proposed MHE design and theory since such a Lyapunov function is generally not known explicitly in practical applications. However, due to Assumption 1 and Proposition 1 some $W_{\delta}$ with $\tilde{P}_{2,\mathrm{x}}$, $\tilde{P}_{2,\theta}$, $\tilde{R}$, $\tilde{Q}\succ 0$ and $\tilde{\eta}\in[0,1)$ is guaranteed to exist. Due to scale invariance of $W_{\delta}$, the weighting matrices $P_{2,\mathrm{x}}$, $P_{2,\theta}$, $R$, $Q\succ 0$ in the objective (17) can be chosen arbitrarily, provided $\eta>\tilde{\eta}$, i.e., $01-\eta>$ 0 is sufficiently small, and the horizon $M$ is long enough (Condition (21) in Theorem 4.2). Compare also a similar discussion in [40, Rem. 1].
@@ 4.3 Motivating Example Revisited
Figure 2: Academic example (Section 3): Absolute state and parameter estimation errors resulting from a standard MHE approach with prior on $\hat{\theta}_{t-M_{t}}$ (2.3) and our proposed MHE approach with prior $\bar{\theta}_{0}$ (4.1).
In the following, we revisit the academic example system (3) discussed in Section 3 and apply the proposed MHE (4.1) to it. We choose the same values for $\bar{\theta}_{0}$, $M$, $\eta$, $Q$, $R$ as for the standard MHE in Section 3, with $P_{2,\mathrm{x}}=P_{\mathrm{x}}$, and $P_{2,\theta}=P_{\theta}$. The corresponding absolute state estimation and parameter errors are shown in Figure 2. For comparison, the estimation results from Section 3 using a standard MHE for state and parameter estimation (2.3) are shown again. As expected from the theoretical analysis (Theorem 4.2), the state estimation error resulting from the proposed MHE is significantly smaller. In particular, the parameter error remains uniformly bounded and the parameter converges to the prior $\bar{\theta}_{0}$ whenever the data is not informative enough.

### v2

@@ 2.3 Standard MHE for State and Parameter Estimation
In the following, we introduce a standard MHE formulation to jointly estimate states and parameters, which is widely applied, see, e.g., [49, 5, 6, 7, 8, 9, 10, 12, 11, 13, 16, 15, 14]. The MHE approach at time step $t$ considers past input data $\{u_{j}\}_{j=t-M_{t}}^{t-1}$ and estimates of past noises and outputs $\{\bar{w}_{j},\bar{y}_{j}\}_{j=t-M_{t}}^{t-1}$ within a window of length $M_{t}=\min\{t,M\}$, with horizon $M\in\mathbb{I}_{\geq 0}$, and the past state and parameter estimates [^3] Typically, this choice is denoted as filtering prior, cf. [40, Chap. 4]., $\hat{x}_{t-M_{t}}$ and $\hat{\theta}_{t-M_{t}}$, respectively. Thereby, the MHE scheme optimizes over a sequence of $M_{t}$ disturbance estimates $\hat{w}_{\cdot|t}=\{\hat{w}_{j|t}\}_{j=t-M_{t}}^{t-1}$ and the initial state and the parameter estimates, $\hat{x}_{t-M_{t}|t}$ and $\hat{\theta}_{t}$, respectively. Using the system model (1), the disturbance sequence, initial state estimate, and parameter estimate define a sequence of state $\{\hat{x}_{j|t}\}_{j=t-M_{t}}^{t}$ and output $\{\hat{y}_{j|t}\}_{j=t-M_{t}}^{t-1}$ estimates. The cost function is chosen as
@@
We denote a (not necessarily unique) minimizer of (7) as $\hat{x}_{t-M_{t}|t}^{*}$, $\hat{w}_{\cdot|t}^{*}$, and $\hat{\theta}_{t}^{*}$, with the corresponding state and output sequences denoted as $\hat{x}_{\cdot|t}^{*}$ and $\hat{y}_{\cdot|t}^{*}$, respectively. The resulting state and parameter estimates at time step $t$ are denoted as $\hat{x}_{t}=\hat{x}_{t|t}^{*}$ and $\hat{\theta}_{t}=\hat{\theta}_{t}^{*}$, respectively. The MHE scheme (7) is applied in receding horizon fashion. Thereby, the current state and parameter estimates are obtained at each time step $t$ by solving the (finite-horizon) MHE problem (7) based on the $M_{t}$ most recent inputs, noise and output estimates, and the prior weightings based on the state and parameter estimates at time step $t-M_{t}$.
Stability of such an MHE scheme can be established using tools from [41] if the augmented state $\begin{bmatrix}x_{t}^{\top}&\theta^{\top}\end{bmatrix}^{\top}$ is detectable, the horizon length $M$ is chosen sufficiently large, and if $\eta\in[0,1)$. Notably, as shown in [50, 51], detectability of the augmented state involves satisfaction of a PE condition, compare Remark 2.2 below for a formal definition of PE. In contrast, the detectability condition considered in this paper (Assumption 1) does not depend on a PE condition, and thus holds uniformly for any sequence of inputs.
Remark 2.2 (Persistency of excitation).
Suppose for simplicity that the system (1) is linearly parameterized
$x_{t+1}$ $=\tilde{f}(x_{t},u_{t})+G(x_{t},u_{t})\theta+E_{\mathrm{x}}w_{t},$ (8a)
$y_{t}$ $=\tilde{h}(x_{t},u_{t},w_{t}),$ (8b)
with $E_{\mathrm{x}}=\begin{bmatrix}I_{n}&0\end{bmatrix}\in\mathbb{R}^{n_{\mathrm{x}}\times n_{\mathrm{w}}}$. Then, we have persistence of excitation [37] over a time window $M$ if there exists a positive constant $c>0$ such that
$PE_{t}\coloneqq\sum_{k=t-M}^{t}G_{k}^{\top}G_{k}\succeq cI,\quad\forall t\in\mathbb{I}_{\geq M},$ (9)
with $G_{k}=G(x_{k},u_{k})$. In case of state measurements, the accuracy of parameter estimates can be bounded based on the signal to noise ratio (SNR), i.e., the minimal singular value of $PE_{t}$ (9) divided by the maximum singular value of the empirical process noise covariance $\Sigma_{t}^{\mathrm{w}}=\sum_{t-M}^{t}E_{\mathrm{x}}w_{t}w_{t}^{\top}E_{\mathrm{x}}^{\top}\in\mathbb{R}^{n_{\mathrm{w}}\times n_{\mathrm{w}}}$. In case this ratio is always sufficiently large, a standard MHE approach for state and parameter estimation can be used to simultaneously estimate states and parameters. However, such an approach can yield arbitrarily bad estimates in case the SNR is small for some time, compare the examples in Sections 3 and 6. Note that in general (9) cannot be verified since only noisy output measurements are available.
@@ 3 Motivating Example
Figure 1: Academic example (Section 3): Absolute state and parameter estimation errors resulting from a standard MHE approach for joint state and parameter estimation (7).
In this section, we provide an academic example to demonstrate that standard MHE formulations for joint state and parameter estimation (Section 2.3) lack fundamental robust stability properties in the absence of PE. The presented example is inspired by [56, Sec. V], which demonstrates parameter drift in adaptive filtering. Consider the system
@@
$y_{t}=$ $\begin{bmatrix}x_{t}+[w_{t}]_{2}\\ \theta x_{t}+[w_{t}]_{3}\end{bmatrix}$, (10b)
with $a=0.99$, $w_{t}]_{1}=0$, $w_{t}]_{2}=-0.1$, and $w_{t}]_{3}=0.1$ for all $t\in\mathbb{I}_{\geq 0}$. Thereby, the constant values of $w_{t}]_{2}$ and $w_{t}]_{3}$ can be interpreted as sensor biases, which often occur in practice. Further consider $x_{0}=1$, $\bar{x}_{0}=1$, $\theta=1$, and $\bar{\theta}_{0}=1.1$, and the MHE formulation (7) with $M=20$, $\eta=0.99$, $Q=I_{3}$, $R=I_{2}$, $P_{\mathrm{x}}=10$, $P_{\theta}=10$, and $\bar{w}_{t}=0$ and $\bar{y}_{t}=y_{t}$ for all $t\in\mathbb{I}_{\geq 0}$. Note that
@@
The input is zero over long periods of time and the state asymptotically converges to zero. Hence, the ratio of the excitation signal compared to the noise becomes very small. As shown in Figure 1, the parameter diverges for such a standard MHE formulation [^4] Note that all MHE approaches applied to this motivating example are implemented in Python using CasADi [57] and the solver IPOPT [58]. for joint state and parameter estimation. Consequently, whenever the output contains informative data (when the input is equal to $1$), the drifted parameter estimate leads to an increasing state estimation error. This phenomena is known as bursting in classical adaptive control [59, 60] and is depicted in the zoom-in in Figure 1. Thus, robust stability of the resulting state and parameter estimate is not guaranteed. Note that this problem equally occurs if an MHE scheme (7) without discounting ($\eta=1$) is applied. In the next section, we show how to overcome this issue by replacing the prior weighting with a regularization on the (constant) prior parameter estimate $\bar{\theta}_{0}$. In particular, we follow ideas from classical adaptive control to prevent ill-conditioned parameter estimation problems, compare our discussion in Section 5.3 below.
@@ 4.1 Moving Horizon Estimation Formulation
Equivalently to the standard MHE formulation for joint state and parameter estimation stated in Section 2.3, our proposed MHE at time step $t$ considers past input data $\{u_{j}\}_{j=t-M_{t}}^{t-1}$ and estimates of past disturbances and outputs $\{\bar{w}_{j},\bar{y}_{j}\}_{j=t-M_{t}}^{t-1}$ in a window of length [^5] The use of the window $M_{t}$ allows us to consider both the initial phase with $t\in\mathbb{I}_{[0,M-1]}$, where we rely on all available measurements, and $t\in\mathbb{I}_{\geq M}$, where we use a fixed horizon $M$. $M_{t}=\min\{t,M\}$, with horizon $M\in\mathbb{I}_{\geq 0}$, and the past state estimate $\hat{x}_{t-M_{t}}$. However, instead of using the past parameter estimate $\hat{\theta}_{t-M_{t}}$, the proposed MHE formulation relies on an available prior $\bar{\theta}_{0}$ of the unknown parameter. The objective of the optimization-based state estimation problem is chosen as
@@
$\min_{\hat{x}_{t-M_{t}|t},\hat{w}_{\cdot|t},\hat{\theta}_{t}}$ $V_{\mathrm{MHE}}^{\mathrm{x}}(\hat{x}_{t-M_{t}|t},\hat{w}_{\cdot|t},\hat{y}_{\cdot|t},\hat{\theta}_{t},t)$ (12a)
@@
We denote a (non-unique) minimizer [^6] We assume that a minimizer to (7) always exists. Existence of such a minimizer can be ensured under mild assumptions, e.g., if $P_{2,\mathrm{x}},P_{2,\theta},Q,R\succ 0$, or the sets $\mathbb{W},\mathbb{Y},\mathbb{X},\Theta$ are compact, cf. [40, App. A.11]. of (12) as $\hat{x}_{t-M_{t}|t}^{*}$, $\hat{w}_{\cdot|t}^{*}$, and $\hat{\theta}_{t}^{*}$, with the corresponding state and output sequences denoted as $\hat{x}_{\cdot|t}^{*}$ and $\hat{y}_{\cdot|t}^{*}$, respectively. The resulting state estimate at time step $t$ is denoted as
@@
The MHE approach (12) is then applied in a receding horizon fashion. Thereby, the state estimate (13) at each time step $t$ is obtained by solving the optimization problem (12) based on the $M_{t}$ most recent inputs, noise and output estimates, and the prior weighting based on the state estimate at time step $t-M_{t}$.
@@ 4.2 Theoretical Analysis
In the following, we establish that the MHE approach (12) is an exponentially $\delta$-IOpS estimator according to Definition 2 provided the horizon $M$ is chosen larger than some lower bound.
@@
with $\rho\in[0,1)$. Then, the MHE estimator (12) is exponentially $\delta$-IOpS according to Definition 2, i.e, (4) holds for all $t\in\mathbb{I}_{\geq 0}$ with (uniform) constants $C_{1},\ C_{2},\ C_{3}>0$, and
$\epsilon=$ $\ \frac{4}{\sqrt{1-\rho}}\sqrt{\frac{\lambda_{\max}(P_{2,\theta})}{\lambda_{\min}(P_{1})}}\sqrt{\rho}^{M}\|\bar{\theta}_{0}-\theta\|,$ (16)
@@
The proof of Theorem 4.3 can be found in Appendix .1. As a consequence of (16), the resulting bound on the state estimation error (4) improves with increasing horizon length $M$ and in case the prior estimate $\bar{\theta}_{0}$ is closer to the true (unknown) parameter $\theta$. Note that choosing the horizon length $M$ poses a trade-off between computational complexity and accuracy, while robustness is preserved provided Condition (15) is satisfied. For $M$ going to infinity, it follows that the MHE approach in (12) is an exponentially $\delta$-IOS estimator according to Definition 2, even for arbitrarily bad prior parameter estimates $\bar{\theta}_{0}$. Consequently, FIE, i.e., Problem (12) with $M_{t}=t$, is exponentially $\delta$-IOS.
@@
Let Assumption 1 hold. Then, the FIE estimator (12) with $M_{t}=t$ is exponentially $\delta$-IOS according to Definition 2.
Proof 4.5.
@@
The above corollary shows that the results from [54, Thm. 9] are recovered.
@@
In case the true parameter is known, i.e., $\bar{\theta}_{0}=\theta$, it follows directly from Theorem 4.3, and in particular (16), that the estimation error (14) resulting from the proposed MHE approach (12) is robustly globally exponentially stable (with $\epsilon=0$) according to [61, Def. 1].
@@
Choosing $P_{2,\mathrm{x}}$, $P_{2,\theta}$, $R$, and $Q$ in (11) based on a $\delta$-IOOS Lyapunov function $W_{\delta}$ (Definition 3) might seem to limit the applicability of the proposed MHE design and theory since such a Lyapunov function is generally not known explicitly in practical applications. However, due to Assumption 1 and Proposition 1 some $W_{\delta}$ with $\tilde{P}_{2,\mathrm{x}}$, $\tilde{P}_{2,\theta}$, $\tilde{R}$, $\tilde{Q}\succ 0$ and $\tilde{\eta}\in[0,1)$ is guaranteed to exist. Due to scale invariance of $W_{\delta}$, the weighting matrices $P_{2,\mathrm{x}}$, $P_{2,\theta}$, $R$, $Q\succ 0$ in the objective (11) can be chosen arbitrarily, provided $\eta>\tilde{\eta}$, i.e., $1-\eta>0$ is sufficiently small, $P_{1}\succ 0$ is chosen sufficiently small, and the horizon $M$ is long enough (Condition (15) in Theorem 4.3), compare also a similar discussion in [41, Rem. 1].
Remark 4.8 (Benefits of online parameter adaptation).
An alternative approach to prevent parameter drift in case of insufficient PE is to not adapt/estimate the parameters online, i.e., to apply the MHE approach (12) with additional constraint $\hat{\theta}_{t}=\bar{\theta}_{0}$. However, such an approach can lead to infeasibilities and persistent state estimation errors even with small noise, compare also the numerical comparison in Section 6.
Remark 4.9 (Slowly varying parameters).
In practical applications the parameter $\theta$ might be slowly varying over time, e.g., due to aging of components. Suppose slow parameter variations, i.e., $\theta_{t}$ to be constant within one horizon length $M$. Then, the proposed MHE approach (12) also ensures $\delta$-IOpS w.r.t. to
$\epsilon=$ $\ \frac{4}{\sqrt{1-\rho}}\sqrt{\frac{\lambda_{\max}(P_{2,\theta})}{\lambda_{\min}(P_{1})}}\sqrt{\rho}^{M}\max_{j\in\mathbb{I}_{[0,t]}}\|\bar{\theta}_{0}-\theta_{j}\|.$
Note that, similar to (16), $\epsilon$ approaches zero for increasing horizon length $M$. Recently, an MHE design to jointly estimate states and time-varying parameters was presented in [62], however, with stability properties depending on the level of PE.
@@ 4.3 Motivating Example Revisited
Figure 2: Academic example (Section 3): Absolute state and parameter estimation errors resulting from a standard MHE approach for joint state and parameter estimation with prior on $\hat{\theta}_{t-M_{t}}$ (7) and our proposed MHE approach with prior $\bar{\theta}_{0}$ (12).
In the following, we revisit the academic example system (10) discussed in Section 3 and apply the proposed MHE approach (12) to it. We choose the same values for $\bar{\theta}_{0}$, $M$, $\eta$, $Q$, $R$ as for the standard MHE in Section 3, with $P_{2,\mathrm{x}}=P_{\mathrm{x}}$, and $P_{2,\theta}=P_{\theta}$. The corresponding absolute state estimation and parameter errors are shown in Figure 2. For comparison, the estimation results from Section 3 using a standard MHE approach for state and parameter estimation (7) are shown again. As expected from the theoretical analysis (Theorem 4.3), the state estimation error resulting from the proposed MHE is significantly smaller. In particular, the parameter error remains uniformly bounded and the parameter converges to the prior $\bar{\theta}_{0}$ whenever the data is not informative enough.